Let X~Bin(50, 0.64). For samples of size 36, calculate P(31.28 < X bar < 33.36).
1
Expert's answer
2020-10-07T16:12:13-0400

"X\sim Bin(50, 0.64)"


"\mu=np=50\cdot0.64=32"

"\sigma=\sqrt{np(1-p)}=\sqrt{50(0.64)(1-0.64)}=2.4\sqrt{2}"

"\bar{X}\sim N(32, (2.4\sqrt{2})^2)"

"n_1=36"


"P(31.28<\bar{X}<33.36)=P(\bar{X}<33.36)-P(\bar{X}<31.28)="

"=P(Z<\dfrac{33.36-32}{(2.4\sqrt{2})/\sqrt{36}})-P(Z<\dfrac{31.28-32}{(2.4\sqrt{2})/\sqrt{36}})\approx"

"\approx P(Z<2.404163)-P(Z<-1.272792)\approx"

"\approx0.991895-0.101546=0.890349"



"P(31.28<\bar{X}<33.36)=0.890349"